OpenStudy (anonymous):

Find sin θ if cot θ = –2 and cos θ < 0. I don't understand this lesson at all.

OpenStudy (anonymous):

can somebody pleaseee help me quickly? The population (in millions) of a certain country can be approximated by the function: P(x)=50*1.02^x where x is the number of years after 2000. Which of the following calculations will tell in what year the population can be expected to reach 100 million? a. ln(2/1.02)+2000 b. ln(2)/ln(1.02)+2000 c. ln(2/1.02) d. ln(2)/ln(1.02)

zepdrix (zepdrix):

Remember your identity for cotangent? Converting it to sines and cosines might give us some information.\[\large \cot \theta=\frac{\cos \theta}{\sin \theta}\]Hmmmm, they told us that cos θ < 0. It's negative. Then how can cot θ be -2? That will only be true if Sin θ is POSITIVE right? Cause then we have a negative cosine over a positive sine, giving us a negative -2 some how. In which quadrant is Cosine negative, and Sine positive?

OpenStudy (anonymous):

is it quadrant II?

zepdrix (zepdrix):

Yesssss, so we're in Quadrant II, and cot θ = -2... hmmm lemme think a sec :d

OpenStudy (anonymous):

they give me options with radicals in the numerator for sin? i don't understand what you mean

zepdrix (zepdrix):

Umm can you give me an idea of what options we have? Then maybe I can figure this out. Do any of them involve sqrt5?

OpenStudy (anonymous):

yes they both do. but they have one with denominator of 5 and the other 2?

zepdrix (zepdrix):

Ok cool I think I've got a handle on this :) Let's see if this makes sense. I'm gonna draw a picture real quick.

OpenStudy (anonymous):

ok.....

zepdrix (zepdrix):

So we know that our angle is somewhere in the second quadrant, so we'll draw a line where our angle is, and then drop a line down to form a triangle with the x-axis. The reason we do this is because we can use our Trig Triangle relationships to label the sides.

OpenStudy (anonymous):

right opp, adj, and hyp?

zepdrix (zepdrix):

Yah! :) So what is Cotangent in relation to the sides? Well we remember that Tangent is OPPOSITE over ADJACENT, yes? So Cotangent will be ADJACENT over OPPOSITE.

zepdrix (zepdrix):

We can think of \[\large \cot \theta = -2 \quad \text{as}\quad \cot \theta = \frac{-2}{1} \quad = \frac{adjacent}{opposite}\]See how we're going to label the sides? :o

OpenStudy (sirm3d):